OpenStudy (anonymous):

Solve for y. 5x – y = 3

OpenStudy (anonymous):

y=5x-3

OpenStudy (anonymous):

5x - y + y = 3 + y 5x = 3+y x = (3+y)/5

OpenStudy (anonymous):

this is hw we find x

OpenStudy (anonymous):

5x - y = 3 y = 5x - 3